OpenStudy (anonymous):

What is an equation for the line that passes through points (8, –3) and (9, 10)?

OpenStudy (ccswims):

first find slope

OpenStudy (ccswims):



OpenStudy (ccswims):

can you do this by yourself?

OpenStudy (luke69603):

the slope is 13 i guess

OpenStudy (ccswims):

ok yea, now to point-slope

OpenStudy (ccswims):

pick any point

OpenStudy (anonymous):

I think I can o-o

OpenStudy (ccswims):

y - y1= m (x -x1)

OpenStudy (anonymous):

^_^ Thanks for the formula.

OpenStudy (anonymous):

O.O I think I got it XD it's (y – 3) = 13(x + 8)... it think.

OpenStudy (anonymous):

*I

OpenStudy (ccswims):

y + 3 = 13x (x - 8)

OpenStudy (ccswims):